# Quenched coalescent for diploid population models with selfing and overlapping generations

### v1

@@ 2.2 The pedigree as important latent variable of the population
Figure 2: (Population process on the right and the corresponding pedigree on the left). A realization of our population process with $N=5$ individuals from time-steps $k=0$ to $5$ in the past (right panel). The corresponding pedigree is shown on the left. Here we consider the Moran model in NWF (25), where $K_{N}=1$ deterministically. The black lines correspond to reproductive relationships while the yellow lines correspond to an individual persisting from one time-step to the next. The yellow lines give rise to overlapping generations.
@@
The population pedigree, or simply the pedigree, refers to an undirected multi-graph $\mathcal{G}_{N}$ with vertex set $N]\times\mathbb{Z}_{+}$, where a vertex $(\ell,k)$ represents the individual with index $\ell$ at the $k$-th time-step. The edges of $\mathcal{G}_{N}$ are between vertices in consecutive time-steps. For each $k\in\mathbb{Z}_{+}$, there are two edges connecting each of the $K^{(k)}_{N}$ children with its parent(s). If the child has two distinct parent (outcrossing), then there is one edge to each parent; if the child has only one parent (selfing), then there are two edges connecting the child to its parent (hence $\mathcal{G}_{N}$ is a multigraph). For each of the remaining $N-K^{(k)}_{N}$ individuals, there is a single edge connecting to itself in the consecutive time-steps.
The importance of the pedigree, as described in greater detail in the section “Previous Work on Pedigrees” of DFBW (24), is that it encapsulates population dynamics that are common to every single locus on the genome. Even loci extremely far apart are coupled by its dynamics. In particular, gene genealogies far apart on the genome are described by conditionally independent realizations of the ancestral process with respect to the pedigree. Gene genealogies, even conditional on the pedigree, are stochastic due to Mendelian randomness. This is demonstrated in Figure 3. Classical coalescent theory implicitly averages over realizations of the pedigree to determine the average gene genealogy. However, this annealing forces distinct loci to be independent, which is a priori unjustifiable due to the stochasticity of the gene genealogies conditional on the pedigree. Indeed, early biological works of BNA (90); WKLR (12) demonstrate how the pedigree can affect the structure of gene genealogies, and the mathematically rigorous works of Tyu (15); DFBW (24); NWF (25); ABFW (25) have begun to show new scaling limits, and the limitations of the annealing of classical coalescent theory.
@@ 3 Ancestral lines and the coalescent process for a sample
Next, we let $n\in\{2,3,\ldots\}$ be the sample size and consider the lineages of $n$ sampled gene copies under our model, when the population size $N$ is larger than $n$. Suppose we sample $n$ distinct gene copies $\{X_{i}^{N}(0)\}_{1\leq i\leq n}\subset\{0,1\}\times[N]$ from $n$ distinct individuals $\{\hat{X}_{i}^{N}(0)\}_{1\leq i\leq n}\subset[N]$, one gene copy from each individual, at time-step 0 (the present). This assumption can be relaxed, see Remark 4.9.
@@ 3.1 The ancestral partition process for a sample of size $n$
Given the pedigree between time-steps k1k-1 and $k$, $\hat{X}_{i}^{N}(k)$ is chosen uniformly among the vertices in time-step $k$ to which $\hat{X}_{i}^{N}(k-1)$ is adjacent.
@@
The size of the state space of the family of coalescing random walks, namely $(\{0,1\}\times[N])^{n}$, is of order $N^{n}$ which is huge when $N$ is large. It is customary to reduce the state space by ignoring the indices of the ancestral individuals and only keep track of the coalescence of sample genealogies as we go backward in time. This is done via partitions of $[n]$, or equivalently, equivalence relations on $[n]$. For $1\leqslant i,j\leqslant n$, $k\in\mathbb{Z}_{+}$, we write $i\sim_{k}j$ if and only if samples $i$ and $j$ descend from the same chromosome $k$ time-steps ago, i.e.,
@@
  • $i$ and $j$ are in the sample block in $\chi^{N,n}(k)$ if and only if $X_{i}(k)=X_{j}(k)$, i.e. the ancestral lines of the $i$-th and the $j$-th samples coalesced $k$ time-steps in the past, and
  • two blocks are in a set together in $\chi^{N,n}(k)$ if and only if $\hat{X}_{i}(k)=\hat{X}_{j}(k)$, i.e. the two extant lineages corresponding to these blocks are in the same individual in the population $k$ time-steps in the past.
@@
the last time-step (up to $k$) at which the ancestral process $\chi^{N,n}$ is totally dispersed (i.e. lies in $\mathcal{E}_{n}$). Note that $S(k)$ is well-defined since by our assumption $\chi^{N,n}(0)\in\mathcal{E}_{n}$. We define $\widetilde{\chi}^{N,n}$ to be the random time-change of $\chi^{N,n}$ as follows:
@@
The $\mathcal{E}_{n}$-valued process $\widetilde{\chi}^{N,n}$ is non-Markovian. However, the random time-change (3.4) not only keeps the process to live in the state space $\mathcal{E}_{n}$ for which our weak convergence method works, but also prevents accumulation of jumps of the process in the limit as the population size $N$ goes to infinity.
@@ 3.2 An annealed scaling limit
Denote by
$\mathfrak{p}_{\xi\eta}^{N,n}:=\mathbb{P}\left(\chi^{N,n}(1)=\eta\mid\chi^{N,n}(0)=\xi\right),$ (3.5)
the one-step transition probabilities for the ancestral process $\chi^{N,n}$. Let $c_{N}=\mathfrak{p}_{\xi_{0}^{2}\mathbf{1}_{2}}^{N,n}$ be the probability that two randomly chosen genes from two distinct individuals coalesce in one time-step in the past. Then
@@
$=$ $\mathbb{E}\left[\frac{(N-K)K}{N^{2}\,(N-1)}\right]\;+\;\frac{1}{\binom{N}{2}}\,\mathbb{E}\!\left[\frac{1}{16}\sum_{i=1}^{N}\bigl(\tilde{V}_{i}^{\,2}-\tilde{V}_{i}\bigr)-\frac{S}{8}\right],$ (3.6)
@@
$d_{N}:=\mathbb{P}\left(\chi^{N,2}(1)=\Big\{\{1\},\{2\}\Big\}\mid\chi^{N,2}(0)=\Big\{(\{1\},\{2\})\Big\}\right)=\mathbb{E}\left[\frac{K_{N}-S_{N}}{N}\right]=\frac{1-\alpha_{N}}{N}\mathbb{E}\left[K_{N}\right].$ (3.8)
@@
Our third and last assumption is on the generator of the ancestral process. Let $F_{\rm hap}:\mathcal{S}_{n}\to\mathcal{E}_{n}$ denote the haploid map defined by
$\{(C_{1},C_{2}),\ldots,(C_{2x-1},C_{2x}),C_{2x+1},\ldots,C_{b}\}\mapsto\{C_{1}\cup C_{2},\ldots,C_{2x-1}\cup C_{2x},C_{2x+1},\ldots,C_{b}\}.$ (3.10)
@@
Define the matrix $\mathit{P}_{N,n}=\left(\mathit{p}_{\xi\eta}^{N}\right)_{\xi,\eta\in\mathcal{E}_{n}}$ by
$\mathit{p}_{\xi\eta}^{N}:=\mathbb{P}\left(F_{\rm triv}\left(\chi^{N,n}(1)\right)=\eta\,\Big|\,\chi^{N,n}(0)=\xi\right)=\sum_{\zeta\in\mathcal{S}_{n}:\,F_{\rm triv}(\zeta)=\eta}\mathfrak{p}_{\xi\zeta}^{N}.$
@@
$Q_{n}:=\lim_{N\to\infty}c_{N}^{-1}\left(\mathit{P}_{N,n}-I\right)$
exists as a real-valued matrix $Q_{n}=(q^{n}(\xi,\eta))_{\xi,\eta\in\mathcal{E}_{n}}$. Furthermore, $\left(Q_{n}\right)_{n\geq 2}$ is a consistent family, which means that for any $m<n$ the pushforward of $Q_{n}$ under $\varpi_{n,m}$ is $Q_{m}$ (i.e. $q^{n}(\xi,\eta)=q^{m}\big(\varpi_{n,m}(\xi),\varpi_{n,m}(\eta)\big)$ for all $\xi,\eta\in\mathcal{E}_{n}$), where $\varpi_{n,m}:\mathcal{E}_{n}\to\mathcal{E}_{m}$ is the projection obtained by restricting the partition to $[m]$.
To describe the annealed limit, we also describe an $n$-$\Xi$-coalescent following (BLS, 18, P.4).
@@
$\Delta:=\{x=(x_{1},x_{2},\ldots)\in[0,1]^{\infty}:x_{1}\geq x_{2}\geq\ldots,\sum_{i=1}^{\infty}x_{i}\leq 1\}.$
An $n$-$\Xi$-coalescent is a continuous-time Markov process $\chi^{n}=\left(\chi^{n}(t)\right)_{t\in\mathbb{R}_{+}}$ taking values in $\mathcal{E}_{n}$, whose transition rates are invariant under permutation and are described as follows: For any $\xi$ in $\mathcal{E}_{n}$ of consisting of $b$ blocks, suppose that $\eta$ is an element of $\mathcal{E}_{n}$ obtained from $\xi$ by keeping $s$ of the blocks of $\xi$ the same, and coalescing the remaining bsb-s blocks into $r$ blocks of sizes $k_{1},k_{2},\ldots,k_{r}.$ The transition rate from $\xi$ to $\eta$ is given by
$\lambda_{b;k_{1},\ldots,k_{r};s}:=\mathbbm{1}_{(r,s)=(1,b-2)}\Xi(\mathbf{0})+\int_{\Delta\setminus\{\mathbf{0}\}}\sum_{l=0}^{s}\sum_{i_{1},\ldots,i_{r+l}\text{ distinct}}^{\infty}\binom{s}{l}\prod_{m=1}^{r}x_{i_{m}}^{k_{m}}\prod_{m=r+1}^{r+l}x_{i_{m}}(1-\left|x\right|)^{s-l}\frac{\Xi(dx)}{\left\langle x,x\right\rangle},$
@@
Informally, an $n$-$\Xi$-coalescent can be described in terms of interval partitions. Take $x\in\Delta\setminus\{\mathbf{0}\}$ at rate $\frac{1}{\left\langle x,x\right\rangle}\Xi$. As $\sum_{i}x_{i}\leq 1$, $x$ decides an interval partition. We throw each block of the coalescent uniformly at random on the unit interval. If they belong to the same element of the partition, and if they are at most $\sum_{i}x_{i}$, then they coalesce. Additionally, each pair of blocks coalesces independently at rate $\Xi(\mathbf{0})$.
@@
We shall show in Lemma Appendix A.1 that, under Assumption 3.3 and a usual condition on the ordered offspring distribution, we have that Assumption 3.6 holds and that for each $n\geq 2$, $Q_{n}$ is the generator of an $n$-$\Xi$-coalescent; i.e. $\lambda_{b;k_{1},\ldots,k_{r};s}=q^{n}(\xi,\eta)$ whenever $\xi,\eta\in\mathcal{E}_{n}$ and $\lambda_{b;k_{1},\ldots,k_{r};s}$ are as in Definition 3.7.
Assumption 3.6 implies that there is a unique finite measure $\Xi$ on $\Delta$ for which $Q_{n}$ is the infinitesimal generator of an $n$-$\Xi$-coalescent governed by $\Xi$ for all $n\geq 2$. This is precisely the content of (Ber, 04, Proposition 3), which shows that the generator $C$ of the coalescence part of an exchangeable fragmentation-coalescence (EFC) process is completely determined by the consistent family $\left(Q_{n}\right)$ of its finite-dimensional restrictions, each $Q_{n}$ encoding the jump rates on $\mathcal{E}_{n}$. Moreover, (Ber, 04, Proposition 3) proves that any such consistent and exchangeable collection of generators $\left(Q_{n}\right)_{n\geq 2}$ arises as the finite-dimensional projections of an EFC process taking values in $\mathcal{E}_{\infty}$. In the purely coalescent case, this implies that any consistent family of $n$-$\Xi$-coalescents can be realized as the projections of an infinite $\mathcal{E}_{\infty}$-valued Markov process with infinitesimal generator $Q$ determined by $\Xi$.
@@
Suppose that Assumptions 3.3, 3.4 and 3.6 hold. Then there is a unique finite measure $\Xi$ on $\Delta$, made precise in Remark 3.8, such that $\bar{\chi}^{N,n}$, as a $\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)$-valued random variable, converges in law to an $n$-$\Xi$-coalescent governed by $\Xi$ as $N\to\infty$.
@@
This result complements that given in BLS (18) by providing the high-selfing analogue of Theorem 1.1 contained therein. That Theorem 3.9 is not extended to the case $d_{N}c_{N}^{-1}\to\infty$ is not due to any particular difficulty with the annealed proof. Rather, the tools for the quenched proof for the the case $d_{N}c_{N}^{-1}\to\infty$ are sufficiently different from those required for the present work.
The proof of Theorem 3.9 is postponed to Section Appendix B, as it will follow from results in Section 4.

### v2

@@ 2.2 The pedigree as important latent variable of the population
Figure 2: (Full single-locus population process on the right and the corresponding pedigree on the left). A realization of our population process with $N=5$ individuals from time-steps $k=0$ to $5$ in the past (right panel). The corresponding pedigree is shown on the left. Here we consider the Moran model in NWF (25), where $K_{N}=1$ deterministically. The black lines correspond to reproductive relationships while the yellow lines correspond to an individual persisting from one time-step to the next. The yellow lines give rise to overlapping generations.
@@
The population pedigree, or simply the pedigree, refers to an undirected multi-graph $\mathcal{G}_{N}$ with vertex set $N]\times\mathbb{Z}_{+}$, where a vertex $(\ell,k)$ represents the individual with index $\ell$ at the $k$-th time-step. The edges of $\mathcal{G}_{N}$ are between vertices in consecutive time-steps. For each $k\in\mathbb{Z}_{+}$, there are two edges connecting each of the $K^{(k)}_{N}$ offspring with its parent(s). If the offspring has two distinct parents (outcrossing), then there is one edge to each parent; if the offspring has only one parent (selfing), then there are two edges connecting the offspring to its parent (hence $\mathcal{G}_{N}$ is a multigraph). For each of the remaining $N-K^{(k)}_{N}$ individuals, there is a single edge connecting to itself in the consecutive time-steps.
The importance of the pedigree, as described in greater detail in the section “Previous Work on Pedigrees” of DFBW (24), is that it encapsulates population dynamics that are common to every single locus on the genome. Even loci extremely far apart are coupled by its dynamics. In particular, gene genealogies far apart on the genome are described by conditionally independent realizations of the ancestral process with respect to the pedigree. Gene genealogies, even conditional on the pedigree, are stochastic due to Mendelian randomness. This is demonstrated in Figure 3. Classical coalescent theory implicitly averages over realizations of the pedigree to determine the average gene genealogy. However, simulations have demonstrated ways in which the pedigree can affect the structure of gene genealogies BNA (90); WKLR (12); WKW (16). Further, mathematically rigorous works have begun to identify new scaling limits of coalescent processes conditional on the pedigree Tyu (15); DFBW (24); NWF (25); ABFW (25).
@@ 3 Ancestral lines and the coalescent process for a sample
Next, we let $n\in\{2,3,\ldots\}$ be the sample size and consider the lineages of $n$ sampled gene copies under our model. We further supposed that these $n$ distinct gene copies come from $n$ distinct individuals, one gene copy from each, so that $n\leq N$ whereas in total there are $2N$ gene copies available in the population. Our analysis of the ancestral process of the sample relies on this assumption that ancestral lineages begin in different individuals. Remark 4.10 shows how our results may be extended to other types of samples, given our focus on high probabilities of selfing ($\alpha_{N}\to 1$). In what follows, we use $X$ to denote labels of gene copies and $\hat{X}$ to denote labels of individuals containing them.
@@ 3.1 The ancestral partition process for a sample of size $n$
Given the pedigree between time-steps $k+1$ and $k$, $\hat{X}_{i}^{N}(k+1)$ is chosen uniformly among the vertices in time-step $k+1$ to which $\hat{X}_{i}^{N}(k)$ is adjacent.
@@
The size of the state space of the family of coalescing random walks, namely $(\{0,1\}\times[N])^{n}$, is of order $N^{n}$. It is customary to reduce the state space by ignoring the indices of the ancestral individuals and only keep track of the coalescence of sample genealogies as we go backward in time. This is done via partitions of $[n]$, or equivalently, equivalence relations on $[n]$. For $1\leqslant i,j\leqslant n$, $k\in\mathbb{Z}_{+}$, we write $i\sim_{k}j$ if and only if samples $i$ and $j$ descend from the same chromosome $k$ time-steps ago, i.e.,
@@
  • $i$ and $j$ are in the same block in $\chi^{N,n}(k)$ if and only if $X_{i}(k)=X_{j}(k)$, i.e. the ancestral lines of the $i$-th and the $j$-th samples coalesced $k$ time-steps in the past, and
  • two blocks are in a set together in $\chi^{N,n}(k)$ if and only if $\hat{X}_{i}(k)=\hat{X}_{j}(k)$, i.e. the two ancestral lineages corresponding to these blocks are in the same individual in the population $k$ time-steps in the past.
@@
in other words the first time-step in the past that the ancestral lineages are not all in distinct individuals. Recall that the sample begins in this totally dispersed state by assumption: $\chi^{N,n}(0)\in\mathcal{E}_{n}$ in the present notation. Thus $S(k)$ is well-defined, as the time for the sample to leave the subset $\mathcal{E}_{n}$ of $\mathcal{S}_{n}$.
We define $\widetilde{\chi}^{N,n}$ to be the random time-change of $\chi^{N,n}$ as follows:
@@
The $\mathcal{E}_{n}$-valued process $\widetilde{\chi}^{N,n}$ is non-Markovian. However, in the limit $N\to\infty$, the random time-change (3.4) not only keeps the process to live in the state space $\mathcal{E}_{n}$ for which our convergence method works, but also prevents any accumulation of jumps of our rescaled process; see Remark 3.3.
@@ 3.2 An annealed scaling limit
Let $c_{N}$ be the probability that two randomly chosen genes from two distinct individuals coalesce in one time-step in the past. Then
@@
$=$ $\mathbb{E}\left[\frac{(N-K)K}{N^{2}\,(N-1)}\right]\;+\;\frac{1}{\binom{N}{2}}\,\mathbb{E}\!\left[\frac{1}{16}\sum_{i=1}^{N}\bigl(\tilde{V}_{i}^{\,2}-\tilde{V}_{i}\bigr)-\frac{\sum_{i=1}^{N}V_{i,i}}{8}\right],$ (3.5)
@@
Remark 3.3.
A main purpose of the random time-change $S$ is to avoid an accumulation of jumps (as $N\to\infty$) that would prevent convergence of the annealed coalescent in the $J_{1}$ topology. Because selfing can coalesce distinct blocks in the same individual on a time-scale much shorter than the coalescence time-scale, in the absence of this random time-change $S$ there will be an accumulation of jumps as $N\to\infty$. Concretely, consider the state $\xi=\{(\{1\},\{2\}),(\{3\},\{4\})\}$. Because we allow for selfing, the process $\left(\chi^{N,n}\left(\lfloor tc_{N}^{-1}\rfloor\right)\right)_{t\in\mathbb{R}_{+}}$, beginning in state $\xi$, could transition to the state $\{(\{1\},\{2\}),\{3,4\}\}$ in one time-step, and then $\{\{1\},\{2\},\{3,4\}\}$ in the next time-step. These two jumps, close together in the $c_{N}$ time-rescaling, would prevent convergence in $J_{1}$.
The issue of accumulation of jumps was dealt with in BLS (18) by considering the complete dispersal map, which views the blocks in the same individual as distinct. This works in BLS (18) because there is no selfing there, and so with probability tending to $1$ the blocks will all disperse before there is any additional overlap of blocks in the same individual.
@@
$d_{N}:=\mathbb{P}\left(\chi^{N,2}(1)=\Big\{\{1\},\{2\}\Big\}\mid\chi^{N,2}(0)=\Big\{(\{1\},\{2\})\Big\}\right)=\frac{1-\alpha_{N}}{N}\mathbb{E}\left[K_{N}\right].$ (3.7)
@@
Let $F_{\rm hap}:\mathcal{S}_{n}\to\mathcal{E}_{n}$ denote the haploid map defined as follows: for
$\xi=\{(C_{1},C_{2}),\ldots,(C_{2x-1},C_{2x}),C_{2x+1},\ldots,C_{b}\}$
we define
$F_{\rm hap}\left(\xi\right):=\{C_{1}\cup C_{2},\ldots,C_{2x-1}\cup C_{2x},C_{2x+1},\ldots,C_{b}\}.$ (3.9)
@@
Our third and last assumption is on the generator of the ancestral process. Let
$\mathfrak{p}_{\xi\eta}^{N,n}:=\mathbb{P}\left(\chi^{N,n}(1)=\eta\mid\chi^{N,n}(0)=\xi\right)$ (3.10)
be the one-step transition probabilities for the ancestral process $\chi^{N,n}$, and define the matrix $\mathit{H}_{N,n}=\left(h_{\xi\eta}^{N}\right)_{\xi,\eta\in\mathcal{E}_{n}}$ by
$h_{\xi\eta}^{N}:=\mathbb{P}\left(F_{\rm hap}\left(\chi^{N,n}(1)\right)=\eta\,\Big|\,\chi^{N,n}(0)=\xi\right)=\sum_{\zeta\in\mathcal{S}_{n}:\,F_{\rm hap}(\zeta)=\eta}\mathfrak{p}_{\xi\zeta}^{N}.$
@@
$Q_{n}:=\lim_{N\to\infty}\frac{1}{c_{N}}\left(\mathit{H}_{N,n}-I\right)$
exists as a real-valued matrix $Q_{n}=(q^{n}(\xi,\eta))_{\xi,\eta\in\mathcal{E}_{n}}$. Furthermore, $\left(Q_{n}\right)_{n\geq 2}$ is a consistent family, meaning that $Q_{n}=Q_{m}\circ\varpi_{n,m}$ for any $2\leq m<n$, where $\varpi_{n,m}:\mathcal{E}_{n}\to\mathcal{E}_{m}$ is the projection obtained by restricting the partition to $[m]$.
To describe the annealed limit, we also describe an $n$-$\Xi$-coalescent (also called $n$-ksi-coalescent governed by $\Xi$ in this paper) following (BLS, 18, P.4).
@@
$\Delta:=\left\{x=(x_{1},x_{2},\ldots)\in[0,1]^{\infty}:x_{1}\geq x_{2}\geq\ldots,\sum_{i=1}^{\infty}x_{i}\leq 1\right\}.$
An $n$-$\Xi$-coalescent is a continuous-time Markov process $\chi^{n}=\left(\chi^{n}(t)\right)_{t\in\mathbb{R}_{+}}$ taking values in $\mathcal{E}_{n}$, whose transition rates are invariant under permutation and are described as follows: For any $\xi$ in $\mathcal{E}_{n}$ consisting of $b$ blocks, suppose that $\eta$ is an element of $\mathcal{E}_{n}$ obtained from $\xi$ by keeping $s$ of the blocks of $\xi$ the same, and coalescing the remaining bsb-s blocks into $r$ blocks of sizes $k_{1},k_{2},\ldots,k_{r}$ respectively. The transition rate from $\xi$ to $\eta$ is given by
$\lambda_{b;k_{1},\ldots,k_{r};s}:=\mathbbm{1}_{(r,s)=(1,b-2)}\Xi(\mathbf{0})+\int_{\Delta\setminus\{\mathbf{0}\}}\sum_{l=0}^{s}\sum_{i_{1},\ldots,i_{r+l}\text{ distinct}}^{\infty}\binom{s}{l}\prod_{m=1}^{r}x_{i_{m}}^{k_{m}}\,\cdot\prod_{w=r+1}^{r+l}x_{i_{w}}\cdot\,(1-\left|x\right|)^{s-l}\frac{\Xi(dx)}{\left\langle x,x\right\rangle},$
@@
Informally, an $n$-$\Xi$-coalescent can be described in terms of interval partitions. Take $x\in\Delta\setminus\{\mathbf{0}\}$ at rate $\frac{1}{\left\langle x,x\right\rangle}\Xi$ and perform an $x$-merger. Additionally, each pair of blocks coalesces independently at rate $\Xi(\mathbf{0})$. In particular, any given fixed pair of lineages coalesce with rate
$\lambda_{2;1;0}=\Xi(\mathbf{0})+\Xi(\Delta\setminus\{\mathbf{0}\})=\Xi(\Delta).$
Writing $x=(x_{1},x_{2},\ldots)$, an $x$-merger can be roughly described as follows: as $\sum_{i}x_{i}\leq 1$, $x$ decides an interval partition. We throw each block of the coalescent uniformly at random on the unit interval. If they belong to the same element of the partition, and if they are at most $\sum_{i}x_{i}$, then they coalesce.
@@
We shall show in Lemma Appendix B.2 that, under Assumption 3.4 and a usual condition on the ordered offspring distribution, we have that Assumption 3.7 holds and that for each $n\geq 2$, $Q_{n}$ is the generator of an $n$-$\Xi$-coalescent; i.e. $\lambda_{b;k_{1},\ldots,k_{r};s}=q^{n}(\xi,\eta)$ whenever $\xi,\eta\in\mathcal{E}_{n}$ and $\lambda_{b;k_{1},\ldots,k_{r};s}$ are as in Definition 3.8.
Remark 3.10.
Assumption 3.7 implies that there is a unique finite measure $\Xi$ on $\Delta$ for which $Q_{n}$ is the infinitesimal generator of an $n$-$\Xi$-coalescent governed by $\Xi$ for all $n\geq 2$. This is the content of (Ber, 04, Proposition 3), which shows that the generator $C$ of the coalescence part of an exchangeable fragmentation-coalescence (EFC) process is completely determined by the consistent family $\left(Q_{n}\right)$ of its finite-dimensional restrictions, each $Q_{n}$ encoding the jump rates on $\mathcal{E}_{n}$. Moreover, (Ber, 04, Proposition 3) proves that any such consistent and exchangeable collection of generators $\left(Q_{n}\right)_{n\geq 2}$ arises as the finite-dimensional projections of an EFC process taking values in $\mathcal{E}_{\infty}$. In the purely coalescent case (i.e. no fragmentation), this implies that any consistent family of $n$-$\Xi$-coalescents can be realized as the projections of an infinite $\mathcal{E}_{\infty}$-valued Markov process with infinitesimal generator $Q$ determined by $\Xi$.
@@
Suppose that Assumptions 3.4, 3.5 and 3.7 hold. Then there is a unique finite measure $\Xi$ on $\Delta$, made precise in Remark 3.10, such that $\bar{\chi}^{N,n}$, as a $\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)$-valued random variable, converges in distribution to an $n$-$\Xi$-coalescent governed by $\Xi$ as $N\to\infty$.
The proof of Theorem 3.11 is postponed to Section Appendix A, as it will follow from results in Section 4.
@@
Theorem 3.11 complements (BLS, 18, Theorem 1.1) by providing the high-selfing analogue. Note that in BLS (18); ABFW (25) where selfing is excluded, the $\Xi$ arising in the annealed limit is such that $\lambda_{2;1;0}=1$. In this paper, however, due to high selfing rate ($\alpha_{N}\to 1$), the measure $\Xi$ in Theorem 3.11 satisfies $\lambda_{2;1;0}=2$, i.e.
$2=\Xi(\mathbf{0})+\Xi(\Delta\setminus\{\mathbf{0}\})=\Xi(\Delta).$ (3.11)
This normalization was also adopted in NWF (25).
Remark 3.13.
That Theorem 3.11 is not extended to the case $d_{N}c_{N}^{-1}\to\infty$ is not due to any particular difficulty with the annealed proof. Rather, the tools for the quenched proof for the the case $d_{N}c_{N}^{-1}\to\infty$ are sufficiently different from those required for the present work. Hence we leave this case to future work. For the simple case of $n=2$ samples from two different individuals, if $\alpha_{N}\to\alpha\in[0,1]$, one can check as in (NWF, 25, Theorem 3.1) that the coalescence time $\tau^{N,2}$ satisfies that $\tau^{N,2}\,c_{N}$ converges in distribution to an exponential random variable with rate $\frac{2}{2-\alpha}$ (i.e. $\lambda_{2;1;0}=\frac{2}{2-\alpha}$). This is consistent with (3.11) when $\alpha=1$.